jabez177 (jabez177):

A mountain bike wheel has a diameter of 26 in. What is the approximate circumference of the wheel? Use pi star times = 3.14. Express your final answer to the nearest hundredth. in. ?

OpenStudy (anonymous):

google circumference of a circle formula

OpenStudy (anonymous):

C=2πr

jabez177 (jabez177):

26 x pi x 2 = 163.28

jimthompson5910 (jim_thompson5910):

that formula works if you use the radius

jabez177 (jabez177):

163.28 / 2 = 81.64

jabez177 (jabez177):

81.64 is my answer?

jimthompson5910 (jim_thompson5910):

diameter = 26 radius = 13 (divide diameter by 2)

jimthompson5910 (jim_thompson5910):

C = pi*d or C = 2*pi*r C = circumference of the circle pi = 3.14 approximately d = diameter r = radius